Use the References to access important values if needed for this questio What concentration of sodium hydroxide is needed to give an aqueous solution with a pH of 11.630? Molarity of sodium hydroxide Submit Answer Try Another Version 8 item attempts remaining 1 pt 1 pt FS FS 2 4
Answer #1


use:
pH = -log [H+]
11.63 = -log [H+]
[H+] = 2.344*10^-12 M

use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(2.344*10^-12)
[OH-] = 4.266*10^-3 M

[NaOH] = [OH-]
= 4.266*10^-3 M
Answer: 4.266*10^-3 M